# SLAM; Point Cloud Registration; LiDAR; Multi-robot SLAM

### v1

@@ 1 Introduction
Moreover, SLAM in perceptually challenging environments, such as underground tunnels, poses difficulties to current systems. A typical SLAM pipeline includes an incremental part, built around the estimation of the motion of the robot between a sensor activation and the subsequent one (aka odometry), as well as a mechanism for integrating such sensor data. Another component of a typical SLAM pipeline is a module to detect loop closures, i.e., the robot passing again at the same place. Lastly, a module for exploiting the loop closures, in order to temper the unavoidable geometric distorsions arising in the incremental phase.
A lot of work has been done with respect to LiDAR odometry estimate in challenging environments, with recent studies mainly focused on the degeneracy of ICP optimization Tuna et al. (2024, 2025), ICP being the long-live state-of-the-art for merging, i.e., registering, point clouds (PCs) from LiDAR activations. However, little has been done w.r.t. the task of place recognition, loop closure and global point cloud registration in the same kind of environments.
@@ 2.1 Dataset description
The simulated environment is based on a high-fidelity mesh reconstructed from real LiDAR scans of the Zentrum am Berg underground infrastructure research facility [^1] https://www.zab.at/en/about-us/the-facility. The data gathered consists of 261 single point clouds, acquired with a RIEGL VZ-400i scanner, providing a ranging accuracy of 5 mm. Since the standard registration of point clouds of a tunnel does not really work well, because the single point clouds do not show much difference since they more or less represent always a tube, the registration have been eased by adding constraints based on the usage of the ad-hoc mounted bireflex targets. These targets, widely used in geomatic surveys in auto-similar conditions like those in underground tunnels, produce a very distinguishable profile in the intensity channel of the point cloud, and this information is then used to help to align the point clouds. The mesh used for the simulations has been obtained from such well-aligned point clouds.
@@
Figure 1: Ground truth trajectories of the four robots, displayed over the mesh of the environment
@@ 2.3 Multi-robot SLAM implementation
To facilitate loop closure detection, which in a multi-robot case means it has to take place also across robots, we compute a global descriptor for each keyframe. Due to the low generalization capabilities of DNN-based descriptors for point clouds, which was demonstrated both by Fontana et. al Fontana et al. (2025) and Seo et al. Seo et al. (2025) on the point cloud registration task, we decided to use the well-known handcrafted Scan Context Kim and Kim (2018) descriptor. For every keyframe from the a robot, we search for the most similar descriptor among the keyframes of a second robot, discarding candidates whose similarity scores fall below a predefined threshold. When a promising match is found, we estimate the relative transformation between the corresponding point clouds using the global registration method KISS-Matcher Lim et al. (2024), a fast and robust approach based on matching FPFH features, and a graph-based outlier rejection scheme, which demonstrates both fast and accurate results.
@@
To ensure robustness in the presence of incorrect loop closures, the main problem that is affecting current multi-robot SLAM systems, we incorporate two complementary techniques: (i) the Pairwise Consistency Measurement (PCM)Mangelson et al. (2018) is used to pre-filter inconsistent inter-robot loop closures before optimization: this method maximizes the consistency between the single robot odometry trajectories and the loops found in the inter robot matches, by representing the consistent measurements as edges of a graph and then finding the maximum clique; and (ii) the Graduated Non-Convexity (GNC) optimizer Yang et al. (2020), which is applied as a final step to improve robustness during the pose graph optimization process: GNC solves non-convex optimization problem by starting with a convex approximation, gradually increasing its non-convexity, and using each solution as the starting point for the next step until the original problem is reached, thus avoiding local minima. As for the GNC optimization, we manually mark odometry edges as inliers, so that they will not be affected by the robust optimization.
To investigate the impact of uninformative data on loop detection and closure, and specifically for our case of “tunnel” structures, we introduce a simple filtering heuristic aimed at excluding keyframes that bring in a small amount of new information. We compute the oriented bounding box, using PCA to compute the orientation of the box Zhou et al. (2018), of each keyframe point cloud and inspect its second-largest dimension. If this dimension falls below a predefined threshold, tuned according to knowledge of the environment, the keyframe is presumed to originate from a tunnel-like structure with limited structural variation and is excluded from the insertion into the graph as well as the subsequent matching and processing process. While this approach is relatively coarse and environment-dependent, it provides an initial means to explore how redundant or uninformative observations influence the performance of multi-robot SLAM.
@@ 3.1 Single robot LiDAR Odometry
The results, for which we use the Absoulute Error Trajectory, i.e. the translation error between the gt poses and the estimated poses, show that KISS-ICP is not resilient against degenerate environments, while Kinematic ICP is robust and can effectively estimate the trajectory, as shown in fig. 2.
Figure 2: Comparison of Absolute Trajectory error for KISS-ICP and Kinematic ICP. On the x axis the 4 trajectories, on the y axis the ATE. The error is represented as a standard boxplot with the median error shown as line in the middle of the box.
@@ 3.2 Multi-robot map merging
Here we present the results of the map merging. Given the distributed approach that we are following, we use the term "map merging" because it integrates the loop closures of each robot altogether with the actual merging across the maps built by each robot. We test the map merging for each robot pair, resulting in 6 different configurations, obtained by varying the robots involved in the pair. We also analyze the impact of using PCM filtering and the use of all keyframes, i.e., those considered by the usual heuristics creating a keyframe every 0.5m motion by a robot, compared with using only filtered keyframes. In tables 2 and 3, we show how the filtering process influences the percentage of outliers present in the final graphs. We can clearly see that our approach eliminates a large percentage of outliers, both wrong loops added because of a wrong place recognition detection, and wrong loops added because the place was correctly recognized but the place recognition algorithm failed. Moreover, in table 4 we show the impact of the reduction in the loop outliers and consider SLAM successful if the overall structure of the final optimized map is consistent with the expected map. Further visual information is provided in Appendix A and Appendix B.
All KF Filtered KF
@@
Table 2: Outlier metrics for robot pairs across All and Filtered keyframes (KF) cases before PCM, expressed as percentage over the total number of loop closures.
All KF Filtered KF
@@
Table 3: Outlier metrics for robot pairs across All and Filtered KF cases after PCM, expressed as percentage over the total number of loop closures.
Robot pair PCM filtering All KF wrong PR Filtered KFs
@@
Table 4: “SLAM all KFs” shows the SLAM success when all keyframes were used for loop closure detection, “SLAM all KFs” shows the SLAM success when filtering out tunnel-like point clouds from the loop closure detection. We define ”success“ if all the robots have a ratio between max ATE and path length less then 1%
Robot pair Percentage 1st robot Percentage 2nd robot
@@
Table 5: Maximum ATE, in meters, of each robot for the experiments with filtered KFs
An interesting fact that can be noticed is that un-informative keyframes influence both the number of outliers in the Place Recognition task, due to global perceptual aliasing, and in the Point Cloud registration task, due to the low geometric structure that makes the alignment fail due to perceptual aliasing in local point cloud patches. By appling the Keyframe filtering we retain only the point clouds that are both globally distinctive from others and also present enough local distinctiveness to allow the correct estimation of the alignment.
The experimental results indicate that relying solely on a robust optimizer is generally not sufficient to ensure reliable SLAM performance, if no common frame is imposed between the robots. In these cases, the optimizer struggles to handle the ambiguities and noise introduced by the unfiltered keyframe set. Nevertheless, its performance shows a noticeable improvement when applied in conjunction with keyframe filtering, suggesting that even a basic level of data pruning can enhance robustness in the optimization phase.
The Pairwise Consistency Maximization (PCM) approach demonstrates strong performance overall and is notably more reliable than the optimizer alone. However, it still falls short in certain challenging configurations. Specifically, when all keyframes are used without any form of filtering, PCM achieves successful results in only two out of six test cases. This indicates that while PCM significantly improves robustness, it is not entirely resilient to the negative effects of unfiltered data.
The most promising results are achieved when PCM is combined with tunnel-based keyframe filtering. This configuration leads to a success rate of five out of six, marking a substantial improvement over other methods. The single failure observed in this setup likely stems from the lack of sufficient high-quality loop closures, possibly due to the overly aggressive exclusion of informative keyframes. Nonetheless, the overall reliability of this approach highlights the importance of combining robust verification mechanisms with intelligent data selection.
An especially noteworthy observation emerges from the configuration denoted as (0,2) + PCM. In this case, the SLAM pipeline fails when tunnel filtering is applied but succeeds when all keyframes are retained. This counterintuitive result underscores a critical limitation of our current keyframe filtering strategy: while intended to improve efficiency and robustness, the filtering method may inadvertently discard keyframes that contain essential spatial information necessary for accurate loop closure. This finding suggests a clear need for more refined and adaptive keyframe selection techniques, capable of preserving structurally significant frames while still reducing computational overhead.
@@ 4 Conclusions
This study presented a practical multi-robot LiDAR SLAM system designed for underground tunnel environments, where traditional localization methods struggle due to severe perceptual aliasing and lack of global positioning. We uncovered new critical challenges that are often overlooked or managed by human intervention or ad-hoc solutions such as having a predefined global reference frame for all robots. In particular, we identified that keyframe selection for loop closures plays a pivotal role in overall system performance, with poor selections leading to unstable or failed optimizations. Furthermore, we observed that certain pairs of point clouds, especially those with minimal geometric variation, are inherently difficult to align reliably, introducing significant errors into the global map. By introducing a simple yet effective keyframe selection strategy, we mitigated some of these issues, achieving consistent results despite the adverse conditions. These findings highlight the need for further research into robust keyframe management and alignment strategies, especially in distributed SLAM systems operating in degenerate environments.
@@ Appendix A Pose graph visualization
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization
@@
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization
@@
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization
@@
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization
@@
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization
@@
(d) Initial g2o graph with filtered KFs
(e) g2o graph with filtered KFs after PCM
(f) g2o graph with filtered KFs after optimization

### v2

@@ 1 Introduction
Moreover, SLAM in perceptually challenging environments, such as underground tunnels, poses difficulties to current systems. A typical SLAM pipeline includes an incremental part, built around the estimation of the motion of the robot between a sensor activation and the subsequent one (aka odometry), as well as a mechanism for integrating such sensor data. Another component of a typical SLAM pipeline is a module to detect loop closures, i.e., the robot passing again at the same place. Lastly, a module for exploiting the loop closures, in order to temper the unavoidable geometric distorsions arising in the incremental phase. Oftentimes, in the case of usage of LiDARs, the incremental phase is called tout-court LiDAR odometry, even though the estimation of the rototranslation between two consecutive activations is just the first of the activities implied, before the fusion of the point clouds
A lot of work has been done with respect to LiDAR odometry estimate in challenging environments, with recent studies mainly focused on the degeneracy of ICP optimization Tuna et al. (2024, 2025), ICP being the long-live state-of-the-art algorithm for merging, i.e., registering, point clouds (PCs) from LiDAR activations. However, little has been done w.r.t. the task of place recognition, loop closure and global point cloud registration in the same kind of environments.
@@ 2.1 Dataset description
The simulated environment is based on a high-fidelity mesh reconstructed from real LiDAR scans of the Zentrum am Berg underground infrastructure research facility [^1] https://www.zab.at/en/about-us/the-facility. The data gathered consists of 261 single point clouds, acquired with a RIEGL VZ-400i scanner, providing a ranging accuracy of 5 mm. Since the standard registration of point clouds of a tunnel does not really work well, because the single point clouds do not show much difference since they more or less represent always a tube, the registration have been eased by adding constraints based on the usage of ad-hoc mounted bireflex targets, show in in fig 1. These targets, widely used in geomatic surveys in auto-similar conditions like those in underground tunnels, produce a very distinguishable profile in the intensity channel of the point cloud, and this information is then used to help to align the point clouds. The mesh used for the simulations has been obtained from such well-aligned point clouds.
Figure 1: Example of a bireflex target, used for aligning the scans to generate the mesh environment
@@
Figure 2: Ground truth trajectories of the four robots, displayed over the mesh of the environment. Starting points shown with colored a dot.
@@ 2.3 Multi-robot SLAM implementation
To facilitate loop closure detection, which in a multi-robot case means it has to take place also across robots, we compute a global descriptor for each keyframe. Due to the low generalization capabilities of DNN-based descriptors for point clouds, which was demonstrated both by Fontana et. al Fontana et al. (2025) and Seo et al. Seo et al. (2025) on the point cloud registration task, we decided to use the well-known handcrafted Scan Context Kim and Kim (2018) descriptor. For every keyframe from a robot, we search for the most similar descriptor among the keyframes of a second robot, discarding candidates whose similarity scores fall below a predefined threshold. When a promising match is found, we estimate the relative transformation between the corresponding point clouds using the global registration method KISS-Matcher Lim et al. (2024), a fast and robust approach based on matching FPFH features, and a graph-based outlier rejection scheme, which demonstrates both fast and accurate results.
@@
To ensure robustness in the presence of incorrect loop closures, the main problem that is affecting current multi-robot SLAM systems, we incorporate two complementary techniques: (i) the Pairwise Consistency Measurement (PCM) Mangelson et al. (2018) is used to pre-filter inconsistent inter-robot loop closures before optimization: this method maximizes the consistency between the single robot odometry trajectories and the loops found in the inter robot matches, by representing the consistent measurements as edges of a graph and then finding the maximum clique; and (ii) the Graduated Non-Convexity (GNC) optimizer Yang et al. (2020), which is applied as a final step to improve robustness during the pose graph optimization process: GNC solves non-convex optimization problem by starting with a convex approximation, gradually increasing its non-convexity, and using each solution as the starting point for the next step until the original problem is reached, thus avoiding local minima. For what concerns the GNC optimization, we mark the odometry edges as inliers, so that they will not be affected by the GNC optimization.
To investigate the impact of uninformative data on loop detection and closure, and specifically for our case of “tunnel” structures, we introduce a simple filtering heuristic aimed at excluding keyframes that bring in a small amount of new information. Our proposed heuristic is based on computing the oriented bounding box of each keyframe point cloud; PCA is used to compute the orientation of the box Zhou et al. (2018). We then inspect the second-largest dimension of the bounding box. In order to detect, and discard, continuous tunnel-like keyframes, we check whether the second largest dimension of the bounding box, likely the width of the tunnel in a tunnel-only scenario, falls below a predefined threshold, tuned according to knowledge of the environment. If the condition is true, the keyframe is presumed to originate from a tunnel-only structure, with limited structural variation like, e.g., an intersection between tunnels, and is excluded from the insertion into the graph as well as from the subsequent matching and processing.
While this approach is relatively coarse and environment-dependent, it provides an initial means to explore how redundant or uninformative loop detection observations influence the performance of multi-robot SLAM, which is the objective of this work.
@@ 3.1 Single robot LiDAR Odometry
The results, for which we use the Absolute Error Trajectory, i.e. the sum of the translation errors between the Ground Truth poses and the estimated poses, show that KISS-ICP is not resilient against degenerate environments, while Kinematic ICP is robust and can effectively estimate the trajectory, as shown in fig. 3.
Figure 3: Comparison of Absolute Trajectory error for KISS-ICP and Kinematic ICP. On the x axis the 4 trajectories, on the y axis the ATE, in meters. The error is represented as a standard boxplot with the median error shown as line in the middle of the box.
@@ 3.2 Multi-robot map merging
Here we present the results of the map merging. Given the distributed approach that we are following, we use the term "map merging" because it integrates the loop closures of each robot altogether with the actual merging across the maps built by each robot. We test the map merging for each robot pair, resulting in 6 different configurations, obtained by varying the robots involved in the pair. We also analyze the impact of using PCM filtering and the use of all keyframes, i.e., those considered by the usual heuristics creating a keyframe every 0.5m motion by a robot, compared with using only tunnel-filtered keyframes. In tables 2 and 3, we show how our proposed filtering process, which we refer to as ”tunnel-filtered KF“ influences the percentage of outliers present in the final graphs. We can clearly see that our approach eliminates a large percentage of outliers, both wrong loops added because of a wrong place recognition detection, and wrong loops added because the place was correctly recognized but the point cloud registration algorithm failed. We consider the PCR failed if the translation error exceeds 1 meter or the rotation error exceeds 15 degrees. Moreover, in table 4 we show the impact of the reduction in the loop outliers and consider SLAM successful if the overall structure of the final optimized map is consistent with the expected map. Further visual information is provided in Appendix A and Appendix B.
All KF Tunnel-filtered KF
@@
Table 2: Percentage of outliers (classified as wrong PR or wrong PCR) among total loop detections for robot pairs, across All and Tunnel-filtered KF cases before PCM
All KF Tunnel-filtered KF
@@
Table 3: Percentage of outliers (classified as wrong PR or wrong PCR) among total loop detections for robot pairs, across All and Tunnel-filtered KF cases after PCM filtering.
Robot pair PCM filtering All KF Tunnel-filtered KFs
@@
Table 4: “All KFs” shows the SLAM successes when all keyframes were used for loop detection, “Tunnel-filtered KFs” shows the SLAM success when filtering out tunnel-like point clouds from the loop detection. We define ”success“ if all the robots have a ratio between max ATE and path length less then 1%
Robot pair ATE 1st robot ATE 2nd robot
@@
Table 5: Maximum ATE, in meters, of each robot for the experiments with tunnel-filtered KFs
An interesting fact that can be noticed is that un-informative keyframes influence both the number of outliers in the Place Recognition task, due to global perceptual aliasing, and in the Point Cloud registration task, due to the low geometric structure that makes the alignment fail due to perceptual aliasing in local point cloud patches. By applying the keyframe filtering we retain only the point clouds that are both globally distinct from others and also present enough local distinctiveness to allow the correct estimation of the alignment.
The experimental results indicate that relying solely on a robust optimizer is generally not sufficient to ensure reliable SLAM performance, if no common frame is imposed between the robots, which is a realistic assumption. In these cases, the optimizer struggles to handle the ambiguities and noise introduced by the tunnel keyframes. On the other hand, its performance shows a noticeable improvement when applied in conjunction with keyframe tunnel-filtering, suggesting that even a basic level of data pruning, like the one proposed, can largely enhance robustness in the optimization phase.
The Pairwise Consistency Maximization (PCM) approach demonstrates strong performance overall and is notably more effective than the optimizer alone. However, it still falls short in certain challenging configurations. Specifically, when all keyframes are used without any form of filtering heuristic, PCM achieves successful results in only two out of six test cases. This indicates that while PCM significantly improves robustness, it is not entirely resilient to the negative effects of tunnel-like unfiltered data.
The most promising results are achieved when PCM is combined with tunnel-based keyframe filtering. This configuration leads to a success rate of five out of six, marking a substantial improvement over other methods. The single failure observed in this setup likely stems from the lack of sufficient high-quality loop closures, possibly due to a too aggressive exclusion of informative keyframes. Nonetheless, the overall reliability of this approach highlights the importance of combining robust verification mechanisms with intelligent data selection.
An especially noteworthy observation emerges from the configuration denoted as (0,2) with PCM filtering in Table 4. In this case, the SLAM pipeline fails when tunnel filtering is applied but succeeds when all keyframes are retained. This counterintuitive result underscores a critical limitation of our current keyframe filtering strategy: while intended to improve efficiency and robustness, the filtering method may inadvertently discard keyframes that contain essential spatial information necessary for accurate loop detection. This finding suggests a clear need for more refined and adaptive keyframe selection techniques, capable of preserving structurally significant frames, while discarding keyframes that are not useful for the loop detection.
@@ 4 Conclusions
This study presented a practical multi-robot LiDAR SLAM system for underground tunnel environments, where traditional localization methods struggle due to severe perceptual aliasing and lack of global positioning. We uncovered new critical challenges that are often overlooked or managed by human intervention or ad-hoc solutions such as having a predefined global reference frame for all robots. In particular, we identified that keyframe selection for loop detection plays a pivotal role in overall system performance, with poor selections leading to unstable or failed optimizations, even with robust optimizers. Furthermore, we observed that certain pairs of point clouds, especially those with minimal geometric variation, are inherently difficult to align reliably, introducing significant errors into the global map. By introducing a simple yet effective keyframe selection strategy, we mitigated some of these issues, achieving consistent results. These findings highlight the need for further research into robust keyframe management and alignment strategies, especially in distributed SLAM systems operating in difficult environments.
@@ Appendix A Pose graph visualization
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization
@@
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization
@@
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization
@@
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization
@@
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization
@@
(d) Initial g2o graph with tunnel-filtered KFs
(e) g2o graph with tunnel-filtered KFs after PCM
(f) g2o graph with tunnel-filtered KFs after optimization